# AgentArch: A Comprehensive Benchmark to Evaluate Agent Architectures in Enterprise

### v1

@@ 3.2 Evaluation Metrics
Our primary evaluation metric is the Acceptable Score, which measures the percentage of records that satisfy all three critical success conditions simultaneously:
$\text{Acceptable}=\frac{|\{r\in R:C(r)\land A(r)\land O(r)\}|}{|R|}\times 100\%$
@@
We define our main acceptability metric in this way because it comes closest to representing success in a real world agentic system. Simply arriving at the expected decision may be insufficient in enterprise deployments. If the model arrives at the right choice, that a leave request should be submitted example, but then does not call the tool to submit the request or notify the employee, the task is not complete. Furthermore, if the model selects all the correct tools and makes the correct final decision but uses the wrong input values for a tool, that could lead to adverse outcomes such as a request being made for the wrong employee or a database being updated with a hallucinated value. For this reason, we consider an agentic flow to be successful if and only if all the expected tools have been called with expected values and the correct final decision is made based on the results of the tool calls.
Tool Choice Evaluation We evaluate tool choice under two conditions:
@@
Tool Arguments Scoring Tool arguments receive percent scoring where $A(r)=1.0$ only when all arguments are correct (compared to ground truth).
Reliability Metrics To assess system reliability, we evaluate each configuration over $k=8$ attempts and compute pass@1—the percent of correct samples over k trials) (DeepSeek-AI et al. 2025) and passˆK—the probability of all k trials succeeding (Yao et al. 2025). The latter metric is particularly important in real world enterprise domains where reliability and consistency is critical.
@@
Additional Metrics We track supplementary behavioral metrics including:
@@ 4.1 Overall Results
The benchmark results reveal that complex enterprise workflows remains a challenging task even for state-of-the-art language models in agentic systems. The simpler requesting time-off task (TO) showed had a peak score of 70.8% (GPT-4.1), while the complex customer routing task (CR) proved significantly more difficult, with scores peaking at only 35.3% (Sonnet 4). Only three models—GPT-4.1, Sonnet 4, and o3-mini—demonstrated meaningful capability on the complex customer request routing task, with most configurations achieving only single-digit scores. Performance varied dramatically across architectural configurations, indicating that no single architecture is optimal for all models. Furthermore, models achieved peak scores on different configurations between use cases suggesting that the optimal agentic configuration is in part dependent on the complexity of the task.
@@
With regards to architecture elements, thinking tools consistently improved performance for non-reasoning models on the simpler time off task. GPT-4.1 improved from 48.5% to 70.8% with thinking tools enabled in single-agent function calling with summarized memory. However, thinking tools provided minimal benefit for o3-mini (55.8% to 56.7%) in its best performing category. Thinking tools showed minimal impact on the complex customer routing use case across all models. Compared to other elements, variation in memory management styles and orchestration strategies (within multi-agent architectures) both had minimal impact on scores.
@@
Function calling generally outperformed ReAct across most models. The notable exception was LLaMA 3.3 70B, which achieved its peak score of 12.2% using single-agent ReAct on the requesting time off task. Multi-agent ReAct consistently underperformed suggesting that ReAct may be better suited towards single-agent systems. Hallucinations for all models (except for GPT-4o) were found exclusively under ReAct settings suggesting that while models may be robust to hallucinations on the function calling schemas they have been trained on, they still have the propensity to hallucinate in ReAct settings. This is particularly relevant on the more complex use case in multi-agent systems. Sonnet 4, for example, showed hallucination rates of 36-36% on ReAct in multi-agent systems but 0% hallucination in every other configuration. See Table 4 for hallucination rates.
@@
While some models achieved their highest overall scores on single-agent architectures, the correct final decision scores for the complex customer routing use case were generally higher for multi-agent architectures. Although Sonnet 4’s highest overall scores were all achieved with single-agent function calling (all scores above 33%), its highest correct final decision scores were for function calling with multi-agent systems (84-87%) while for single agent it was in the 72-76% range. GPT-4.1 showed similar results as it achieved its peak overall score in both use cases with single agent function calling; however, on the customer routing task it achieved correct final decision rates of 97-99% with multi-agent function calling and only 79-86% with single agent function calling. GPT-4.1-mini also achieved significantly higher correct final decision rates with multi-agent function calling than with single-agent function calling on both use cases. This suggests that while some models are more likely to choose the correct tool or tool arguments on single agent systems, multi-agent systems may be more effective for selecting the correct final decision after all tool calls have been made. See Table 5 for more details.
@@
Two key patterns emerged from the consistency analysis shown in Table 3. Sonnet 4 and GPT-4.1 showed the most robust performance with relatively low coefficients of variation (32.1% and 27.0% respectively for the simpler task), making them the most reliable choices across different agentic configurations. Conversely, o3-mini exhibited extreme sensitivity to architectural choices (CV = 143.7%).
@@
Reliability Analysis. Passˆk scores across all models and agentic configurations peak at 0.0634 indicating only a 6.34% chance of executing the workflow correctly in all 8 trials. In enterprise settings, where reliability and consistency are essential, pass@1 and passˆk results reveal a fundamental gap between the promise of agentic LLMs and their real-world performance. See appendix A.4 for full scores.
4.2 Model Specific Results
TO CR TO CR TO CR TO CR TO CR TO CR
Orch-Open React Complete Enabled 71.0 75.0 14.6 30.2 10.6 4.0 15.6 5.8 1.7 0.0 87.9 68.0
Complete Disabled 70.8 73.2 13.3 26.2 10.6 3.8 13.8 3.8 1.0 0.0 86.5 65.2
Function Calling Complete Enabled 73.3 97.5 69.0 64.0 83.3 89.5 83.3 92.2 0.0 0.0 75.2 85.2
Complete Disabled 66.0 98.5 64.8 64.8 77.7 89.0 83.5 91.8 0.0 0.0 64.2 87.5
Summarized Enabled 74.8 99.2 69.8 64.8 79.8 86.8 83.8 89.0 0.0 0.0 76.0 86.8
Summarized Disabled 69.4 98.5 68.1 64.0 74.6 84.5 81.9 91.0 0.0 0.0 67.3 86.5
Orch-Isolated React Complete Enabled 71.9 83.5 11.5 33.0 7.3 4.5 15.8 4.2 0.4 0.0 88.3 63.7
Complete Disabled 72.7 82.0 13.8 31.2 8.5 5.0 14.4 4.0 0.8 0.0 88.5 70.5
Function Calling Complete Enabled 71.7 99.2 66.7 65.0 85.2 92.2 84.2 94.2 0.0 0.0 76.5 84.5
Complete Disabled 72.1 97.8 68.1 64.8 79.2 92.5 82.7 91.5 0.0 0.0 64.2 85.8
Summarized Enabled 79.2 97.8 67.3 64.5 83.8 90.0 82.9 92.5 0.0 0.0 74.0 86.8
Summarized Disabled 75.4 97.8 64.4 66.5 75.0 92.8 82.9 91.0 0.0 0.0 66.9 84.0
Single Agent React Complete Enabled 73.5 57.5 41.7 32.5 59.8 47.5 52.7 58.5 14.6 0.0 74.4 71.5
Complete Disabled 76.7 58.2 26.5 29.0 50.4 49.0 38.5 47.0 21.2 0.0 72.3 70.5
Function Calling Complete Enabled 72.3 86.5 64.2 66.2 59.4 23.5 83.8 18.0 0.0 0.0 76.7 75.5
Complete Disabled 69.4 79.2 62.1 60.5 42.9 21.5 84.8 18.0 17.9 0.0 44.8 72.5
Summarized Enabled 74.4 85.0 61.9 65.0 59.0 23.8 82.3 14.2 8.5 0.0 76.5 76.0
Summarized Disabled 70.4 80.8 58.8 61.8 46.0 22.8 83.3 14.2 0.0 0.0 44.6 74.0
Table 5: Pass@1 Correct Final Decision Rate. Bolded scores indicate highest score for a model on each use case.
@@ 4.2 Model Specific Results
Sonnet 4 achieved the highest scores on the complex customer routing task with a peak score of 35.3% using single-agent function calling. Sonnet 4 performed strongly on the requesting time off task (68.5% using orchestrator-isolated function calling). Notably, thinking tools had a substantial impact on the simpler task but minimal effect on the complex customer routing task. Time off scores improved from 64.4% to 68.5% with thinking tools, while customer routing performance remained relatively flat across configurations (33.3%-35.3% range).
@@
GPT-4.1 achieved the highest overall score on the simpler use case - 70.8% using single-agent function calling with summarized memory and thinking tools. The model showed sensitivity to thinking tools in single-agent configurations, improving from 48.5% to 70.8% when enabled. However, thinking tools provided minimal benefit in multi-agent setups (54.4% vs 58.8%). Memory type preferences varied by orchestration: single-agent performance was nearly identical with complete (68.%1) versus summarized (70.8%) memory, while orchestrated configurations slightly favored complete memory.
@@
o3-mini demonstrated extreme architectural sensitivity, achieving its peak score of 56.7% on the requesting time off use case with single-agent function calling but dropping to as low as 1.3% with orchestrated ReAct. Comparing enabled/disabled configurations showed negligble differences (e.g., 55.8% vs 56.7% in single-agent function calling). The model achieved its highest customer routing score of 22.3% using orchestrator-isolated function calling, substantially outperforming its single-agent CR performance (1.5%-2.0% range). This suggests that o3-mini is not inherently suited to single-agent vs multi-agent architecture but rather that its performance depends heavily on the complexity of the task.
GPT-4o and GPT-4.1-mini exhibited similar patterns, both peaking with orchestrator-isolated function calling configurations (53.5% and 67.1% TO respectively) using complete memory and thinking tools. Both models benefited substantially from thinking tool on simple tasks. GPT-4o improved from 26.9% to 39.0% in single-agent function calling with thinking tools, but they had minimal impact in orchestrated scenarios. GPT-4o performed extremely poorly with ReAct. GPT-4.1-mini, however, achieved higher scores on ReAct (41.5% and 51.2%) than on function calling (19.8-35.4%) within single-agent set ups on the time off use case. This finding illustrates the importance of testing agentic architectural elements in conjunction with each other since GPT-4.1-mini performed drastically differently on ReAct in multi-agent vs single-agent systems.
LLaMA 3.3 70B showed poor overall performance scoring 0 on most configurations. However, it achieved its highest score of 12.2% on the time off use case using single-agent ReAct without thinking tools, making it the only model to perform best with ReAct. LLaMA’s poor performance is likely due to its tool repetition rate which averaged 76.0% on the time off use case. Results on the customer routing case were almost entirely 0 due to persistent difficulties with instruction following and the function-calling format.
@@ 5 Discussion
We now use the above findings to make concrete recommendations for designing agentic AI systems in enterprise.
Avoiding Multi-Agent ReAct. One of the few universal trends that this benchmark shows is that models do not perform their best on Multi-Agent ReAct. In our set up, we used ReAct both for the orchestrator to select agents and for the agents to select tools. It is possible that using ReAct for just one instead of both of these components might improve outcomes but further work would be needed to test this.
Multi-Agent Systems for Final Decisions. While some models performed better overall on single-agent systems, we find that multi-agent systems are generally more effective at arriving at the correct final decision. For enterprise use cases where arriving at a final decision is more important than exact tool selection, we recommend the use of multi-agent systems.
@@
Using Thinking Tools. Our results indicate that for non-reasoning models, thinking tools can be a useful way for enabling more careful analysis on the information collected over the course of an agentic workflow. This is particularly true in contexts where the model must make a calculated decision based on the information it has collected (as is the case with the requesting time off use case). Thinking tools, however, can also add latency so we recommend developers consider carefully if their use cases require the extra reasoning abilities that these tools provide.
Model Selection. While large state-of-the-art models (in our benchmark GPT-4.1 and Claude Sonnet 4) perform the best across architectures, our benchmark shows that on simpler tasks, it is possible to achieve strong performances with smaller models if the appropriate architecture is chosen. For example, while GPT-4.1-mini’s scores on the time off use case dip as low as 4.6%, it’s peak acceptable score of 67.1% is comparable to Sonnet 4’s peak score of 68.5%.
@@ 6 Limitations
Our benchmark has several important limitations. We evaluate only two enterprise use cases with 60 samples each, which may not generalize across all industries and organizational contexts. We design one use case to represent a simple workflow and another as a more complex task, but a wider range of use cases would improve understanding of how task complexity affects performance on various architectures. Another key limitation is our model selection which is limited to six models, with only one open-source model (LLaMA 3.3 70B) and one reasoning-specialized model (o3-mini). Inclusion of more reasoning models in particular would provide more conclusive results on whether or not the results seen for o3-mini can be generalized to reasoning models as a whole. Finally, our use of deterministic tool responses enables controlled evaluation but may not reflect the variability of production enterprise systems.
7 Future Work
Several directions remain for expanding this benchmark to better capture enterprise agentic system requirements.
Multimodal Enterprise Integration. Our current evaluation is limited to text-based tool inputs and outputs, while real enterprise environments frequently involve multimodal data processing including document analysis, image recognition, and file processing workflows. Future work should expand the benchmark to include tools that accept and return images, PDFs, and other file formats commonly encountered in enterprise systems.
@@ 7 Future Work
Multi-Agent ReAct. Our benchmark shows poor performances on multi-agent ReAct across models but does not fully illuminate the reason for this poor performance. Future work focusing specifically on Multi-Agent ReAct and testing ReAct for orchestration and tool calling independently in multi-agent systems would provide more insights on whether ReAct can be effectively utilized in multi-agent systems.
8 Conclusion
We introduce AgentArch, a benchmark for evaluating LLMs on a comprehensive agentic architectures on enterprise use cases. We test 18 configurations (single agent vs multi-agent, ReAct vs function calling, integration of thinking tools, and memory management styles) and find that even state-of-the-art LLMs struggle to maintain reliable performance on complex enterprise workflows in any set up. We also observe model-specific architectural preferences across LLMs suggesting that there is no universally optimal agentic architecture. Furthermore, even the best performing models do not demonstrate reliability over multiple trials showing a fundamental gap between the potential of agentic systems and their effectiveness in practice. We hope these findings provide actionable insights that enterprise developers can use to improve their model selection and architecture design of agentic systems.
Appendix A Appendix
A.1 Use Cases and Set Up

### v2

@@ 3.2 Evaluation Metrics
In enterprise agentic systems, reaching the correct conclusion is necessary but insufficient. Consider a leave request scenario: even if the model correctly determines that leave should be approved, the task fails if the model doesn’t invoke the approval API, calls it with incorrect parameters (e.g., wrong employee ID), or omits required notifications. Real-world agentic deployment requires that the agentic system (1) selects the right tools to complete all required steps, (2) invokes them with correct arguments, and (3) reaches the correct final outcome.
As such, our primary evaluation metric is the Acceptable Score, which measures the percentage of records that satisfy all three critical success conditions simultaneously:
$\textbf{Acceptable}=\frac{|\{r\in R:C(r)\land A(r)\land O(r)\}|}{|R|}\times 100\%$
@@
Tool Choice Evaluation. We evaluate tool choice under two conditions:
@@
Tool Arguments Scoring. Tool arguments receive percent scoring where $A(r)=1.0$ only when all arguments are correct (compared to ground truth).
Reliability Metrics. To assess system reliability, we evaluate each configuration over $k=8$ attempts and compute pass@1—the percent of correct samples over $k$ trials. deepseekai2025deepseekr1incentivizingreasoningcapability and passˆK—the probability of all $k$ trials succeeding taubench:25. The latter metric is particularly important in real-world enterprise domains where reliability and consistency are critical.
@@
Additional Metrics We track supplementary behavioral metrics, including:
@@ 4.1 Overall Results
Figure 4: Pass@1 Hallucination Rate ($\downarrow$). Heatmap cells show the percentage of trials (lower is better, darker indicates higher error) in which an agent hallucinated non-existent entities or schema, e.g., selecting a tool or agent not present in the registry or inventing parameters not defined by a tool schema. Rates are computed at Pass@1 over $k=8$ attempts per model–configuration on both tasks (TO, CR). Hallucinations tend to concentrate in ReAct settings, especially in multi-agent orchestration.
@@
Complex enterprise workflows remain challenging for current agentic systems. On the simpler time-off task (TO), peak performance reaches 70.8% (GPT-4.1), while the more complex customer routing task (CR) peaks at 35.3% (Sonnet 4). Only GPT-4.1, Sonnet 4, and o3-mini show meaningful capability on CR; most other models score in the single digits. Performance varies markedly across architectures, with no universally best configuration. Models also peak on different configurations across the two tasks, indicating that optimal design depends on task complexity.
@@
With regards to architecture elements, thinking tools consistently improved performance for non-reasoning models on the simpler time off task. GPT-4.1 improved from 48.5% to 70.8% with thinking tools enabled in single agent function calling with summarized memory. However, thinking tools provided minimal benefit for o3-mini (55.8% to 56.7%) in its best performing category. Thinking tools showed minimal impact on the complex customer routing use case across all models. Compared to other elements, variation in memory management styles and orchestration strategies (within multi-agent architectures) had minimal impact on scores.
@@
Function calling generally outperformed ReAct across most models. The notable exception was LLaMA 3.3 70B, which achieved its peak score of 12.2% using single agent ReAct on the requesting time off task. Multi-agent ReAct consistently underperformed indicating that ReAct may be better suited towards single agent systems. Hallucinations for all models (except for GPT-4o) were found exclusively under ReAct settings suggesting that while models may be robust to hallucinations on the function calling schemas they have been trained on, they still have the propensity to hallucinate in ReAct settings. This is particularly relevant on the more complex use case in multi-agent systems. Sonnet 4, for example, showed hallucination rates of 36-36% on ReAct in multi-agent systems but 0% hallucination in every other configuration. See Figure 4 for hallucination rates.
@@
While some models achieved their highest overall scores on single agent architectures, the correct final decision scores for the complex customer routing use case were generally higher for multi-agent architectures. Although Sonnet 4’s highest overall scores were all achieved with single agent function calling (all scores above 33%), its highest correct final decision scores were for function calling with multi-agent systems (84-87%), while for single agent, it was in the 72-76% range. GPT-4.1 showed similar results as it achieved its peak overall score in both use cases with single agent function calling; however, on the customer routing task, it achieved correct final decision rates of 97-99% with multi-agent function calling and only 79-86% with single agent function calling. GPT-4.1-mini also achieved significantly higher correct final decision rates with multi-agent function calling than with single agent function calling on both use cases. This suggests that while some models are more likely to choose the correct tool or tool arguments on single agent systems, multi-agent systems may be more effective for selecting the correct final decision after all tool calls have been made. See Figure 5 for more details.
@@
Two key patterns emerged from the consistency analysis shown in Table 2. Sonnet 4 and GPT-4.1 showed the most robust performance with relatively low coefficients of variation (32.1% and 27.0%, respectively, for the simpler task), making them the most reliable choices across different agentic configurations. Conversely, o3-mini exhibited extreme sensitivity to architectural choices (CV = 143.7%).
@@
Reliability Analysis. Passˆk scores across all models and agentic configurations peak at 0.0634, indicating only a 6.34% chance of executing the workflow correctly in all 8 trials. In enterprise settings, where reliability and consistency are essential, pass@1 and passˆk results reveal a fundamental gap between the promise of agentic LLMs and their real-world performance. See Appendix A.4 for full scores.
4.2 Model-Specific Results
Figure 5: Pass@1 Correct Final Decision Rate. Heatmap cells report the probability (higher/ darker is better) that a model–configuration returns the correct ground-truth outcome (approve/deny, route/escalate, etc.). Computed at Pass@1 over $k=8$ attempts on both tasks (TO, CR). Together with Fig. 5 (Acceptable), this highlights a trade-off between decision accuracy and end-to-end execution reliability.
@@ 4.2 Model-Specific Results
Sonnet 4 achieved the highest scores on the complex CR task with a peak score of 35.3% using single agent function calling. Sonnet 4 performed strongly on the TO task (68.5% using orchestrator-isolated function calling). Notably, thinking tools had a substantial impact on the simpler task but a minimal effect on the CR task. TO scores improved from 64.4% to 68.5% with thinking tools, while CR performance remained relatively flat across configurations (33.3%-35.3% range).
@@
GPT-4.1 achieved the highest overall score on the simpler use case - 70.8% using single agent function calling with summarized memory and thinking tools. The model showed sensitivity to thinking tools in single agent configurations, improving from 48.5% to 70.8% when enabled. However, thinking tools provided minimal benefit in multi-agent setups (54.4% vs 58.8%). Memory type preferences varied by orchestration: single agent performance was nearly identical with complete (68.%1) versus summarized (70.8%) memory, while orchestrated configurations slightly favored complete memory.
@@
o3-mini demonstrated extreme architectural sensitivity, achieving 56.7% on time off requests with single agent function calling but dropping to 1.3% with orchestrated ReAct. Thinking tools showed negligible differences (55.8% vs 56.7%). The model achieved its highest CR score of 22.3% using orchestrator-isolated function calling, substantially outperforming single agent performance (1.5%-2.0%). This suggests o3-mini’s performance depends heavily on task complexity rather than inherent suitability to single versus multi-agent architecture.
GPT-4o and GPT-4.1-mini peaked with orchestrator-isolated function calling (53.5% and 67.1% TO) using full memory and thinking tools. Both gained from thinking tools on simple tasks—GPT-4o rose from 26.9% to 39.0% in single-agent setups, though gains faded under orchestration. GPT-4o struggled with ReAct, while GPT-4.1-mini scored higher on ReAct (41.5–51.2%) than function calling (19.8–35.4%) in single-agent mode, underscoring how agentic elements interact differently across architectures.
LLaMA 3.3 70B performed poorly overall, scoring 0 on most configurations. Its best result was 12.2% on the TO task with single-agent ReAct (no thinking tools)—the only model to perform best with ReAct. Scores on customer routing were near zero due to persistent instruction-following issues.
@@ 5 Discussion
We now use the above findings to make concrete recommendations for designing agentic AI systems in an enterprise.
Avoiding Multi-Agent ReAct. One of the few universal trends that this benchmark shows is that models do not perform their best on Multi-Agent ReAct. In our setup, we used ReAct both for the orchestrator to select agents and for the agents to select tools. It is possible that using ReAct for just one instead of both of these components might improve outcomes, but further work would be needed to test this.
Multi-Agent Systems for Final Decisions. While some models performed better overall on single agent systems, we find that multi-agent systems are generally more effective at arriving at the correct final decision. For enterprise use cases where arriving at a final decision is more important than exact tool selection, we recommend the use of multi-agent systems.
@@
Using Thinking Tools. Our results indicate that for non-reasoning models, thinking tools can be a useful way for enabling more careful analysis of information collected over the course of an agentic workflow. This is particularly true in contexts where the model must make a calculated decision based on the information it has collected (as is the case with the requesting time off use case). Thinking tools, however, also add latency, so we recommend developers consider carefully if their use cases require the extra reasoning abilities that these tools provide.
Model Selection. While large models (in our benchmark GPT-4.1 and Claude Sonnet 4) perform the best across architectures, our benchmark shows that on simpler tasks, it is possible to achieve strong performances with smaller models if the appropriate architecture is chosen. For example, while GPT-4.1-mini’s scores on the time off use case dip as low as 4.6%, its peak acceptable score of 67.1% is comparable to Sonnet 4’s peak score of 68.5%.
6 Conclusion
We introduce AgentArch, a benchmark for evaluating LLMs on a comprehensive agentic architectures on enterprise use cases. We test 18 configurations (single agent vs multi-agent, ReAct vs function calling, integration of thinking tools, and memory management styles) and find that even state-of-the-art LLMs struggle to maintain reliable performance on complex enterprise workflows in any setup. We also observe model-specific architectural preferences across LLMs, suggesting that there is no universally optimal agentic architecture. Furthermore, even the best performing models do not demonstrate reliability over multiple trials showing a fundamental gap between the potential of agentic systems and their effectiveness in practice. We hope these findings provide actionable insights that enterprise developers can use to improve their model selection and architecture design of agentic systems.
@@ 7 Limitations
Our benchmark has several limitations.
Use Case Diversity and Size. We evaluate only two enterprise use cases with 60 samples each, which may not generalize across industries or contexts. One use case represents a simple workflow and the other a complex task, but more varied cases would better reveal how task complexity affects architectural performance.
Model selection. Due to cost constraints, our experiments are currently limited to six models total, with only one open-source and one reasoning model. Including more reasoning models would clarify whether o3-mini’s results generalize to reasoning models as a class.
Multimodal Enterprise Integration. Our current evaluation is limited to text-based tool inputs and outputs, while real enterprise environments frequently involve multimodal data processing, including document analysis, image recognition, and file processing workflows. Future work should expand the benchmark to include tools that accept and return images, PDFs, and other file formats commonly encountered in enterprise systems.
@@
Multi-Agent ReAct. Our benchmark shows poor performance on multi-agent ReAct across models, but does not fully illuminate the reason for this poor performance. Future work focusing specifically on Multi-Agent ReAct and testing ReAct for orchestration and tool calling independently in multi-agent systems would provide more insights on whether ReAct can be effectively utilized in multi-agent systems.
Sampling Parameters. All experiments use temperature=0 for reproducibility, which may not reflect production deployments where some stochasticity could be beneficial. The interaction between sampling parameters and architectural choices remains unexplored.
Success Metrics. Our "Acceptable Score" requires perfect alignment on all three dimensions (tools, arguments, outcome). This strict evaluation may not capture scenarios where partial success provides business value. Additionally, we do not measure efficiency metrics (latency, cost, token usage) which are critical for enterprise deployment decisions.
8 Acknowledgments
We are grateful to Sai Rajeswar Mudumba for his careful reading of the manuscript and helpful suggestions and guidance that improved the final version of this work.
Appendix A Use Cases and Set Up